Find the general solution and the second solution of,

x2y"-3xy'+4y=0,given y1(x)=x2

1
Expert's answer
2021-07-19T17:31:02-0400

Answer:-

Given that "y_1(x)=x^2"  is one solution.

"\implies y'_1=2x\\
\implies y''_1=2"

Given D.E "x^2y"-3xy'+4y=0"

Now  "x^2y_1''-3xy'_1+4y_1=x^2(2)-3x(2x)+4x^2=0"

That is. "y_1(x)" satisfies the given DE

So "y_1(x)" s a solution of the D.E.


Let the second solution be "y_2(x)=u(x)x^2"

"\implies y_2'=u'x^2+2ux\\\implies y''_2=u''x^2+2u'x+2[u'x+u]"

or "y_2''=u''x^2+4u'x+2u"


Substitute these into the differential equation to get

"x^2(u''x^2+4u'x+2u)-3x(u'x^2+2ux)+4ux^2=0"

"\implies u''x^4+u'x^3=0\\
\implies u''x+u'=0\ \ \ \ \ \ \ \ \ \ \ \ \ (\because x >0) \\
\implies \frac{u''}{u'}=-\frac{1}{x}"

on integrating we get

"lnu'=-lnx+c\\
\ \ \ \ \ \ \ \ =ln(\frac{1}{x})+c"


By taking c=0 we get

"lnu' =ln(\frac{1}{x})\\
\implies u'=\frac{1}{x}\\
\implies u=lnx"


So second solution "y_2(x)=(lnx)x^2"

"W(x)=\begin{vmatrix}
 x^2 & x^2lnx\\
 2x & 2xlnx+x
\end{vmatrix}"

Now wronskian 

"=2x^3lnx+x^3-2x^3"

"=x^3
\\\implies \boxed{W(x)\ne0} \ \ \ \ \ \ (\because x >0)"

Hence "y_1(x) and Y_2(x)" form a fundamental set of solutions.

So general solution is given by

"\boxed{y=c_1x^2+c_2x^2lnx}"